# Parameter-Efficient Fine-Tuning of State Space Models

### v1

@@ Abstract
Deep State Space Models (SSMs), such as Mamba (Gu & Dao, 2024), have emerged as powerful tools for language modeling, offering high performance with efficient inference and linear scaling in sequence length. However, the application of parameter-efficient fine-tuning (PEFT) methods to SSM-based models remains largely unexplored. This paper aims to systematically study two key questions: (i) How do existing PEFT methods perform on SSM-based models? (ii) Which modules are most effective for fine-tuning? We conduct an empirical benchmark of four basic PEFT methods on SSM-based models. Our findings reveal that prompt-based methods (e.g., prefix-tuning) are no longer effective, an empirical result further supported by theoretical analysis. In contrast, LoRA remains effective for SSM-based models. We further investigate the optimal application of LoRA within these models, demonstrating both theoretically and experimentally that applying LoRA to linear projection matrices without modifying SSM modules yields the best results, as LoRA is not effective at tuning SSM modules. To further improve performance, we introduce LoRA with Selective Dimension tuning (SDLoRA), which selectively updates certain channels and states on SSM modules while applying LoRA to linear projection matrices. Extensive experimental results show that this approach outperforms standard LoRA.
@@ 1 Introduction
Over the past two years, Large Language Models (LLMs) such as ChatGPT (Achiam et al., 2023; Brown et al., 2020) have achieved groundbreaking performance and are now widely used in daily life. Many models use the Transformer architecture (Vaswani et al., 2017), with its attention mechanism essential in predicting subsequent tokens based on context. Each token computes attention scores with every preceding one, selectively focusing only on the most relevant during processing. This, however, creates quadratic time complexity, posing challenges when dealing with long sequences. In response, various alternative architectures like Linear Attention (Katharopoulos et al., 2020), RWKV (Peng et al., 2023), RetNet (Sun et al., 2023), and Mamba (Gu & Dao, 2024) have been developed to operate with subquadratic time complexity.
As the most popular subquadratic-time architecture currently serving as an alternative to Transformers, SSMs  (Gu et al., 2021, 2022b, 2022a; Gu & Dao, 2024) achieve efficient training and inference. SSMs are closely related to linear RNNs, which maintain a hidden state to encapsulate the information of previous tokens. When a new input token is introduced, the prediction of the next token involves only operations on this hidden state and the new token, which enhances inference efficiency. To overcome the limitation of RNNs, which cannot be trained in parallel, S4 (Gu et al., 2022b, a) leverages its linearity, enabling it to adopt a convolutional form during training, facilitating parallel computation. Consequently, SSMs are highly efficient and have demonstrated success in numerous long-sequence tasks (Gu et al., 2022b, a). Recently, a new series of SSM models, Mamba (Mamba-I (Gu & Dao, 2024) and Mamba-II (Dao & Gu, 2024)), have achieved Transformer-level performance in language modeling. Although numerous SSM-based models exist, we primarily focus on the deep S4 model and Mamba. The deep S4 model, serving as the foundational architecture, readily extends its properties to other variants, while Mamba has emerged as one of the most popular SSM-based models.
Consequently, we expect fine-tuning these pretrained SSMs for downstream tasks will become a crucial problem in the near future. While fine-tuning the entire model is expensive and inefficient, numerous Parameter-Efficient Fine-Tuning (PEFT) methods (Houlsby et al., 2019; Hu et al., 2021; He et al., 2021; Li & Liang, 2021; Lester et al., 2021; Zaken et al., 2022; Liu et al., 2021, 2022) have been developed for efficient adaptation under resource constraints. Notably, most popular PEFT methods fall into two categories: (i) prompt-based tuning, which involves modifying the input sequence (Lester et al., 2021) or tuning the sequence at each layer (Li & Liang, 2021); and (ii) parameter-based tuning, which directly updates the model parameters, such as LoRA (Hu et al., 2021), which modifies the weight matrices, and BitFit (Zaken et al., 2022), which updates only the bias terms.
Despite the success that existing PEFT methods have achieved in adapting Transformer-based models, their efficacy in adapting SSM-based models remains largely unexplored, leaving many interesting questions open. For instance, are existing popular PEFT methods still effective for SSM-based models? If they are applicable, what is the optimal way to apply these methods to SSM-based models, and which parameters should be updated? If not, can we develop variants specifically tailored for SSMs that perform better? To answer these questions, to the best of our knowledge, we conduct the first comprehensive study of PEFT on SSM-based models, both theoretically and empirically.
To the best of our knowledge, we are the first to benchmark existing PEFT methods on SSM-based models. Through extensive experiments, we demonstrate that (Finding 1) prompt-based PEFT methods are no longer effective for SSM-based models, and (Finding 2) LoRA remains effective on SSM-based models. Meanwhile, the two major components of SSM-based models are the SSM module, which functions analogously to attention in Transformers, and linear projection matrices, which are similar to feed-forward layers. We next investigate which part of the model is more effective for applying PEFT. We empirically find that (Finding 3) applying LoRA to linear projection matrices without modifying the SSM module is already effective, while the most effective linear projection matrices differ depending on the dataset. Notably, Findings 1 and 3 are supported by our theoretical analysis. While LoRA is not effective for tuning SSM modules, theoretically, tuning additional SSM modules increases expressivity. Finally, we analyze the architecture of SSM-based models using the theoretical framework of Giannou et al. (2023) and Zeng & Lee (2024). We show that, in addition to applying LoRA to linear projection matrices, Selectively updating the channel and state Dimensions of SSM modules further enhances performance. We dub this method as SDLoRA, the first PEFT method tailored for SSM-based models. Through extensive experiments, we observe that (Finding 4) SDLoRA outperforms LoRA alone in fine-tuning SSM-based models.
@@ 2 Related Works
State Space Models (SSMs).
Despite of the remarkable performance of SSMs on certain tasks of sequence modeling, SSMs still showed worse performance than Transformers on language modeling. Fu et al. (2022) transitioned from synthetic language modeling tasks to real language modeling tasks with SSMs. They proposed H3, which is inspired by Linear Attention (Katharopoulos et al., 2020), introducing both diagonal SSM and shift SSM. Recently, Mamba (Gu & Dao, 2024; Dao & Gu, 2024) escaped from linear time invariance (LTI) modeling by introducing input-dependent terms and achieved better performance than Transformer on language modeling. Furthermore, several hybrid models (Lieber et al., 2024; Park et al., 2024) tried to exploit the advantages of both SSMs and Transformers.
Parameter-Efficient Fine-Tuning (PEFT).
Due to the increase in model size, PEFT methods have gained increasing popularity as they achieve good performance while being much more efficient compared to full model updating (Houlsby et al., 2019; Hu et al., 2021; He et al., 2021; Li & Liang, 2021; Lester et al., 2021; Zaken et al., 2022; Liu et al., 2021, 2022). Most of the existing popular PEFT methods fall into two categories: (i) prompt-based methods (Li & Liang, 2021; Lester et al., 2021; Liu et al., 2021, 2022), and (ii) parameter tuning methods (Donahue et al., 2014; Yosinski et al., 2014; Hu et al., 2021; Zaken et al., 2022). Common prompt-based methods include prompt tuning (Lester et al., 2021), and prefix-tuning (Li & Liang, 2021). Prompt tuning prepends a sequence of learnable virtual tokens, which are continuous vectors. Prefix-tuning further expands on Prompt tuning by prepending tokens across the model’s depth, making it more powerful. Therefore, our analysis of prompt-based methods’ limitations will focus on prefix-tuning, with the findings also applicable to the other prompt-based methods. Conversely, parameter tuning methods, which originated from traditional transfer learning practices, typically involve freezing the initial layers and only tuning the last few layers (Donahue et al., 2014; Yosinski et al., 2014). In recent years, more effective and innovative parameter tuning approaches have emerged (Hu et al., 2021; Zaken et al., 2022). The widely used Low-Rank Adaptation (LoRA) updates a subset of parameters (e.g., attention layers of a Transformer) in a low-rank manner. BitFit (Zaken et al., 2022), focuses on tuning only the bias terms of a pretrained model. In Sec. A, we provide a more detailed description of these baseline methods.
Numerous efforts have been made to theoretically understand existing PEFT methods. For prompt-based methods, Wang et al. (2023b), Petrov et al. (2024), and Oymak et al. (2023) have theoretically analyzed the effectiveness and limitations of prompt tuning and prefix-tuning for Transformer-based models. For LoRA, Zeng & Lee (2024) explored its expressive power by demonstrating that even a randomly initialized model can be adapted to match any smaller target model using LoRA. Some of our theoretical analysis draws upon the framework established by Zeng & Lee (2024). Jang et al. (2024) conducted a theoretical exploration of LoRA within the neural tangent kernel (NTK) regime.
3 Preliminaries of State Space Models
Scalar-input Scalar-output SSM.
@@ 3 Preliminaries of State Space Models
To adapt SSMs for deep learning, the continuous parameters $({\bm{A}},{\bm{B}})$ are transformed into discrete counterparts $(\overline{{\bm{A}}},\overline{{\bm{B}}})$ using a learnable step size $\Delta\in{\mathbb{R}}$. An example of a discretization rule is the zero-order hold, which defines $\overline{{\bm{A}}}=\exp(\Delta{\bm{A}}),\overline{{\bm{B}}}=(\Delta{\bm{A}})^{-1}(\exp(\Delta{\bm{A}})-{\bm{I}})\cdot\Delta{\bm{B}}$.
The discrete-time SSM is formulated as (2). For efficient and parallelizable training, the output ${\bm{y}}$ of a length-$N$ input ${\bm{x}}$ in the discrete-time SSM can be computed with a long convolution, as detailed in (3). This convolution operation can be efficiently computed in the frequency domain with FFT.
Vector-input Vector-output SSM.
Many deep learning tasks, such as language modeling, often use multi-channel inputs. When the input and output are vectors, denoted as $\bm{x},\bm{y}\in{\mathbb{R}}^{D}$, separate SSMs are used for each of the $D$ input channels. As such, a superscript $(d)$ is introduced to indicate parameters specific to each channel when necessary. This notation may be omitted for simplicity.
@@
S4 introduced by Gu et al. (2022b) represents one of the earliest applications of SSMs in deep learning. It features a diagonal structure for the state matrix ${\bm{A}}$, a design theoretically validated by Gu et al. (2022b) and practically implemented through its subsequent variants, DSS (Gupta et al., 2022) and S4D (Gu et al., 2022a).
Deep S4 Layer.
@@
$\bm{y}_{t}=\operatorname{ReLU}({\bm{W}}\cdot\operatorname{S4}_{t}(\bm{x}_{1},\ldots,\bm{x}_{t})+{\bm{\beta}}+{\bm{u}}\otimes\bm{x}_{t}),$ (4)
where ${\bm{W}}\in{\mathbb{R}}^{D\times D}$ and ${\bm{\beta}}\in{\mathbb{R}}^{D}$ represent the linear projection matrix and bias, respectively, and ${\bm{u}}\in\mathbb{R}^{D}$ is the coefficient of the residual connection. Note that in a deep S4 layer, the trainable parameters are SSM parameters $({\bm{A}}^{(d)},{\bm{B}}^{(d)},{\bm{C}}^{(d)},\Delta^{(d)})$ across $D$ channels with ${\bm{A}}^{(d)}$ being diagonal and the parameters $({\bm{W}},{\bm{\beta}})$ for the linear layer and ${\bm{u}}$ for the residual connection.
@@
A key property of all SSMs mentioned above is linear time invariance (LTI), where model dynamics remain constant over time. However, LTI models face significant limitations: their constant dynamics fail to selectively extract relevant information from the context or influence the hidden state in an input-dependent manner. The S6 model, proposed by Gu & Dao (2024), addresses these limitations by making its parameters input-dependent.
In particular, at each time step $t$, given the input $\bm{x}_{t}\in{\mathbb{R}}^{D}$, they introduce input-dependency to step size $\bm{\Delta}_{t}=(\Delta^{(1)}_{t},\ldots,\Delta^{(D)}_{t})^{\top}\in{\mathbb{R}}^{D}$, input transition vectors ${\bm{B}}_{t}\in{\mathbb{R}}^{H\times 1}$ and the output mapping vectors ${\bm{C}}_{t}\in{\mathbb{R}}^{1\times H}$ via linear projection:
$\bm{\Delta}_{t}=\operatorname{softplus}(\bm{W}_{\bm{\Delta}}\bm{x}_{t}+{\bm{\beta}}_{\bm{\Delta}}),\quad{\bm{B}}_{t}={\bm{W}}_{{\bm{B}}}\bm{x}_{t},\quad{\bm{C}}_{t}={\bm{W}}_{{\bm{C}}}\bm{x}_{t},$ (5)

### v3

# Kevin Galim    Wonjun Kang    Yuchen Zeng    Hyung Il Koo    Kangwook Lee

@@ Abstract
Deep State Space Models (SSMs), such as Mamba (Gu & Dao, 2024), have become powerful tools for language modeling, offering high performance and linear scalability with sequence length. However, the application of parameter-efficient fine-tuning (PEFT) methods to SSM-based models remains largely underexplored. We start by investigating two fundamental questions on existing PEFT methods: (i) How do they perform on SSM-based models? (ii) Which parameters should they target for optimal results? Our analysis shows that LoRA and its variants consistently outperform all other PEFT methods. While LoRA is effective for linear projection matrices, it fails on SSM modules—yet still outperforms other methods applicable to SSMs, indicating their limitations. This underscores the need for a specialized SSM tuning approach. To address this, we propose Sparse Dimension Tuning (SDT), a PEFT method tailored for SSM modules. Combining SDT for SSMs with LoRA for linear projection matrices, we achieve state-of-the-art performance across extensive experiments.
Machine Learning, ICML
[…]
@@ 1 Introduction
In the past few years, Large Language Models (LLMs) such as ChatGPT (Achiam et al., 2023; Brown et al., 2020) have achieved groundbreaking performance and are now widely used in daily life. While many models rely on the Transformer architecture (Vaswani et al., 2017), its quadratic time complexity due to the attention mechanism poses challenges for long sequences. To address this, alternative architectures such as linear attention (Katharopoulos et al., 2020), RWKV (Peng et al., 2023), RetNet (Sun et al., 2023), and Mamba (Gu & Dao, 2024) have been developed, offering subquadratic time complexity. Efficient attention alternatives often rely on State Space Models (SSMs) or their variants (Gu et al., 2021, 2022b, 2022a; Gu & Dao, 2024), which are akin to linear RNNs, maintaining hidden states of fixed size for sequential processing. S4 (Gu et al., 2022b, a) overcomes RNNs’ parallel training limitations by constraining parameter structures, enabling a convolutional form for efficient parallel computation. S6 (Gu & Dao, 2024) improves this with input-dependent parameters, enabling selective focus on relevant information per token. Building on S6 with linear projection matrices (analogous to the Feed-Forward Networks in Transformer layers), Mamba-I (Gu & Dao, 2024) emerged as a prominent SSM-based model. Mamba-I was later extended to Mamba-II (Dao & Gu, 2024), with both models achieving Transformer-level performance in language modeling and gaining widespread recognition.
As SSMs gain popularity, performing parameter-efficient fine-tuning (PEFT) on pretrained models for downstream tasks is crucial, since full fine-tuning is costly and inefficient. Numerous PEFT methods (Houlsby et al., 2019; Hu et al., 2021; He et al., 2021; Li & Liang, 2021; Lester et al., 2021; Zaken et al., 2022; Liu et al., 2021, 2022; Houlsby et al., 2019) have been developed, achieving notable success on Transformer models. The most popular PEFT methods fall into three categories: (i) input-injection methods, which add sequences to the model’s main input (Lester et al., 2021) or prepend tokens to the intermediate inputs at each layer (Li & Liang, 2021); (ii) architecture-enhancement methods, which adjust the model architecture. For example, Houlsby et al. (2019) added layers between Transformer layers, while Additional-scan (Yoshimura et al., 2025) expands state dimensions in the SSM module; (iii) weight-tuning methods, which directly modify existing model weights. Notable weight-tuning approaches include BitFit (Zaken et al., 2022), which updates only bias terms, and LoRA (Hu et al., 2021), which modifies weight matrices through low-rank updates, along with its variants such as DoRA (Liu et al., 2024) and LoRA+ (Hayou et al., 2024). For simplicity, we denote LoRA and its variants as LoRA.
Figure 1: A visual guide to PEFT methods in SSM-based models: benchmarking and innovation. We compare various existing PEFT approaches on SSM-based models, demonstrating that LoRA applied to linear projection matrices outperforms all other methods. However, extending LoRA to SSM modules fails to yield further improvements. To address this, we propose Sparse Dimension Tuning (SDT), which achieves state-of-the-art performance on SSM-based models when combined with LoRA for linear projection matrices.
Despite the success that existing PEFT methods have achieved in adapting Transformer-based models, their efficacy in adapting SSM-based models remains largely underexplored, leaving many interesting questions open.
  1. 1.
    Do existing popular PEFT methods remain effective for SSM-based models?
  2. 2.
    If applicable, what is the optimal way to integrate these methods into SSM-based models, and which parameters should be updated?
  3. 3.
    If not, can we design specialized variants tailored to SSMs that yield superior performance?
Our main contributions to address these questions are:
  • Comprehensive Benchmarking of PEFT Methods. We benchmark six widely used PEFT methods across three categories on diverse tasks, including natural language understanding, generation, and computer vision. We evaluate these methods on both SSM-based models (i.e., Mamba) and a hybrid model (i.e., Jamba (Lieber et al., 2025)), which consists of both Transformer layers and Mamba layers. Our results show that LoRA consistently outperforms all other PEFT methods on both SSM-based and hybrid models. However, its effectiveness is limited to linear projection matrices, as further tuning of SSM modules does not improve performance. Notably, other methods applicable to SSM modules perform worse than LoRA, further underscoring the need for a specialized approach to tuning SSM modules.
  • Introducing Sparse Dimension Tuning (SDT) for SSM Modules. To develop an effective method for tuning SSM modules, we conduct a theoretical analysis to understand the roles of different parameters. This analysis motivates the Sparse Dimension Tuning and Pruning (SDT-P) method, which improves efficiency by freezing and pruning certain channel and state dimensions while training only the remaining ones. We establish theoretical guarantees for its effectiveness in SSM-based models when combined with LoRA applied to linear projection matrices. We then simplify SDT-P into Sparse Dimension Tuning (SDT) by omitting explicit pruning, as pruned dimensions can be considered equivalent to training dimensions set to zero. SDT selectively updates channels and fine-tunes specific dimensions within them, as illustrated in Fig. 1.
  • Demonstrating Effectiveness of SDT. Through extensive experiments, we demonstrate that integrating SDT into SSM-based models, combined with applying LoRA to their linear projection matrices, achieves state-of-the-art fine-tuning performance.
The roadmap of our paper is illustrated in Fig. 1. Our code is available at https://github.com/furiosa-ai/ssm-peft.
@@ 2 Related Works
Concurrent Works of PEFT on SSMs.
Several concurrent studies (Halloran et al., 2024; Yoshimura et al., 2025; Kang et al., 2025) have investigated PEFT methods for SSM-based models. Halloran et al. (2024) studied both in-context learning and parameter-efficient fine-tuning, with an orthogonal focus on analyzing Mamba’s stability under mixed-precision training using Lyapunov exponents. Kang et al. (2025) introduced state-based PEFT methods and proposed State-offset Tuning, solely focusing fine-tuning Mamba’s S6 blocks. Yoshimura et al. (2025) benchmarked multiple PEFT approaches—including established methods and a new method called Additional-scan (which adds a trainable state dimension to the SSM module), plus partial tuning (fine-tuning only a subset of parameters)—and introduced MambaPEFT through PEFT search strategies. While Yoshimura et al. (2025) solely focused on Mamba-I, providing an in-depth study of that particular architecture, our work investigates a broader class of SSM-based models including deep S4, Mamba-I, Jamba in the main body, as well as Mamba-II presented in Sec. C.2 and E.2, aiming to offer general insights on how to effectively tune SSMs rather than focusing on a single variant.
Sparse Tuning.
Several studies have explored sparse parameter selection in fine-tuning (Song et al., 2024) and skill localization (Panigrahi et al., 2023). Song et al. (2024) showed that sparse tuning is an effective PEFT method, linking the low intrinsic dimensionality of pre-trained models to the proportion of parameters needing updates. They propose selecting optimal fine-tuning parameters based on gradient magnitudes. We enable sparse tuning for SSM by applying sparsity across entire dimensions (channel and state) rather than specific neurons. Panigrahi et al. (2023) focused on identifying neurons responsible for specific downstream tasks by fully fine-tuning the model and computing neuron masks to minimize task loss. While effective for skill localization, this method is computationally expensive and not optimized for parameter-efficient fine-tuning.
In Sec. A, we provide a more detailed discussion of related work on SSMs and PEFT.
3 Preliminaries
3.1 State Space Models
Discrete-Time SSMs.
@@ 3.1 State Space Models
To handle discrete inputs, the continuous parameters $({\bm{A}},{\bm{B}})$ are discretized into $(\overline{{\bm{A}}},\overline{{\bm{B}}})$ using a learnable step size $\Delta\in{\mathbb{R}}$. A common discretization rule, the zero-order hold, defines $\overline{{\bm{A}}}=\exp(\Delta{\bm{A}}),\overline{{\bm{B}}}=(\Delta{\bm{A}})^{-1}(\exp(\Delta{\bm{A}})-{\bm{I}})\cdot\Delta{\bm{B}}$. The discrete-time SSM, given in (2), enables efficient inference via long convolution described in (3). For multi-channel inputs $\bm{x},\bm{y}\in{\mathbb{R}}^{D}$, separate SSMs are used per channel, with a superscript $(d)$ indicating channel-specific parameters when needed.
@@
$\bm{y}_{t}=\operatorname{ReLU}({\bm{W}}\cdot\operatorname{S4}_{t}(\bm{x}_{1},\ldots,\bm{x}_{t})+\bm{\beta}+{\bm{u}}\odot\bm{x}_{t}),$ (4)
where ${\bm{W}}\in{\mathbb{R}}^{D\times D}$ and $\bm{\beta}\in{\mathbb{R}}^{D}$ represent the linear projection matrix and bias, respectively, and ${\bm{u}}\in{\mathbb{R}}^{D}$ is the coefficient of the residual connection. Trainable parameters include SSM parameters $({\bm{A}}^{(d)},{\bm{B}}^{(d)},{\bm{C}}^{(d)},\Delta^{(d)})$ across $D$ channels with ${\bm{A}}^{(d)}$ being diagonal, as well as linear layer $({\bm{W}},\bm{\beta})$ and residual connection ${\bm{u}}$.
@@
All SSMs mentioned above exhibit linear time invariance (LTI), meaning their dynamics remain constant over time. A key limitation of LTI SSMs is their fixed dynamics, hindering selective context extraction and input-dependent state transitions. S6 (Gu & Dao, 2024) addresses this by making parameters input-dependent. At each time step $t$, given input $\bm{x}_{t}\in{\mathbb{R}}^{D}$, S6 introduces input-dependent step sizes $\bm{\Delta}_{t}=(\Delta^{(1)}_{t},\ldots,\Delta^{(D)}_{t})^{\top}\in{\mathbb{R}}^{D}$, input transition vectors ${\bm{B}}_{t}\in{\mathbb{R}}^{H\times 1}$ and output mapping vectors ${\bm{C}}_{t}\in{\mathbb{R}}^{1\times H}$ via linear projection:
$\bm{\Delta}_{t}=\operatorname{softplus}(\bm{W}_{\bm{\Delta}}\bm{x}_{t}+\bm{\beta}_{\bm{\Delta}}),\quad{\bm{B}}_{t}={\bm{W}}_{{\bm{B}}}\bm{x}_{t},\quad{\bm{C}}_{t}={\bm{W}}_{{\bm{C}}}\bm{x}_{t},$